You invested $9,000 in a savings deposit 7 quarters ago and it has grown to $9683 today. What nominal rate of annual interest (compounding quarterly) did you earn? (expressed as a percentage to two decimal places
Answer by VFBundy(438)

9683 = 9000 * (1 + r/4)^7

9683/9000 = (1 + r/4)^7

1.07589 = (1 + r/4)^7

(1.07589)^(1/7) = 1 + r/4

1.01050 = 1 + r/4

0.01050 = r/4

4(0.01050) = r

r = 0.0420

The annual interest rate was 4.20%.